OpenStudy (anonymous):

Wendy is creating a large soup based on a recipe. Her recipe calls for 3/4 of a pint of milk, but she only has 5/8 of a pint of milk, so she can only make a portion of her soup. Wendy can only make of her recipe with the amount of milk that she has.

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

if right

OpenStudy (mertsj):

\[\frac{\frac{5}{8}}{\frac{3}{4}}\]

OpenStudy (anonymous):

i cant read that its really small

OpenStudy (mertsj):

Use control + and make it bigger

OpenStudy (anonymous):

whats the answer

OpenStudy (mertsj):

Whatever you get when you simplify that complex fraction.

OpenStudy (anonymous):

5/6